OpenStudy (anonymous):

help ! :D (Sinx/2 + cosx/2) squared =?

OpenStudy (anonymous):

(Let us actually square it as follows a^2 + 2ab + b^2

OpenStudy (anonymous):

sin^2 + 2sinx/2 cos x/2 + cos^2 x/2

OpenStudy (lgbasallote):

\[\huge (\frac{\sin x}{2} + \frac{\cos x}{2})^2\] you know that\[(a + b)^2 \implies a^2 + 2ab + b^2\] so this one is just \[\Large \implies (\frac{\sin x}{2})^2 + 2(\frac{\sin x}{2} \times \frac{\cos x}{2}) + (\frac{\cos x}{2})^2\] so if you simplify that you get \[\implies \frac{\sin^2 x}{4} + 2\left (\frac{\sin x \cos x}{4} \right) + \frac{\cos ^2 x}{4}\] are you getting the idea?

OpenStudy (anonymous):

Camjan the sum of \[\sin^2 \frac{ x }{ 2 } + \cos^2 \frac{x}{2} = Very Easy\]

OpenStudy (anonymous):

and\[2 \sin \frac{ x }{ 2 } \cos \frac{ x }{ 2 } = \sin 2*\frac{ x }{ 2 } = \sin x\]

OpenStudy (anonymous):

Alltogether the CORRECT SOLUTION is\[1 + \sin x\]

OpenStudy (anonymous):

\[(Sin x/2 + \cos x/2)^2\]

OpenStudy (anonymous):

aand the correct gratitude for correct solution is a M...

OpenStudy (anonymous):

\[\sin^2 x/2 + \cos^2 x/2 + 2\sin x/2 * \cos x/2\]

OpenStudy (anonymous):

\[\left( \sin \frac{ x }{ 2 } + \cos \frac{ x }{2}\right)^{2} = \] maybe for you its easy but for me its not that easy thats why i asked you. i just wanna see the step by step solution tnx :)

OpenStudy (anonymous):

Camjan read my solution and USE the fact that\[\sin^2 \frac{ x }{ 2 } + \cos^2 \frac{ x }{ 2 } = 1\]

OpenStudy (anonymous):

1 + 2sin x/2 * cosx/2 = 1 + sin 2 * x/2 = 1 + sinx

OpenStudy (anonymous):

This is what I meant by "very Easy" Right hand side

OpenStudy (anonymous):

i cant understand what our teacher taught us she's so fast .. pls help me guys my exam is tomorrow

OpenStudy (anonymous):

@camjan Did u understand..

OpenStudy (anonymous):

oh ok :) tnx

OpenStudy (anonymous):

Mr Yahoo seems to re-type EVERY LINE I WROTE TO YOU ( but in bad writing :(

OpenStudy (anonymous):

And by the way - a medal is due, and closing the question

OpenStudy (anonymous):

@Mikael hahaha.. its ok

OpenStudy (anonymous):

oh...Sorry....@Mikael i didnt see..that

OpenStudy (anonymous):

Any way u deserve a Medal